When we arrived to MDRS, the door to the Green Hab was locked and all plants were dead. The door got unlocked the same day and Shannon Rupert did substantive works to put things in order, remove dry plants and prepare pots for new crops.
I have finished today what Shannon started: I have collected trash and removed the resting dry plants as well as I put things in order and brushed the floor. I have also analysed the seeds that are available for planting. Most of the pots contain soil and none of them contains plants (there are still dry roots left though in some of the pots). Approximately a half of them is ready to be used for new plants.
Each morning I turn the fan on and I turn it off in the evening. I also partially open and close a sort of the window built into the door. The fan works well. The days are very sunny, so it gets very hot inside (the temperature outside was 26 C degrees; I still need to find a thermometer to measure the temperature inside).
The next step is to start seeds as well as bring some plants from the field and try to grow them inside the Green Hab. I also need to examine how to regulate the temperature, lightening and moisture of the soil in a way it fits the seeds.
Best wishes, Karolina Zawieska Green Hab Officer, Crew 176
EVA #1 approaching to conglomerate Dacota Formation, 1,5 km from habitat and the study of geology map to identify place of collected samples. These Cretaceous Dacota Sandstone and Conglomerates originated due to very strong flow of water. This formation is comparable with images from Curiosity rover from surface of Martian Gale Crater. The rock consists with 1-2 cm pebbles.
This morning we woke up even earlier than before, i.e. at 7:15 am. This was because we needed to wear holter monitors and make them work. We also did morning sport exercises that are part of holter testing. The training started at 8:50 am and lasted for 35 min. We are planning to repeat the training every day.
After the training we had a quick breakfast and the two of us, including myself, got ready for the EVA. The EVA was scheduled for 10:00 am, however, due to the communication issues, we left the hab at 10:20 am (for detailed information see the EVA Report). Due to the heat we experienced due to the EVA, both of us felt tired and took a nap after coming back to the hab.
In the meantime, the Crew Engineer worked on a small mobile robot for approximately 2 hours and the Green Hab Officer spent around 1 hour working in the Green Hab. We had a late lunch at 4 pm. The strategy we have invented for flushing the toilet continues to work.
In the afternoon, the two of the crew members, namely Crew Engineer & Green Hab Officer took a rental car and went to Salt Lake City to pick the last member of the crew up and collect the delayed package (the latter includes such things as the rover Ares 2, tools and equipment). They will be back tomorrow (15th of March).